### v2

# revtex4-1Repair the float Journal of UAP Studies        Society for UAP Studies

Abstract
Claims of the retrieval of crashed craft and vehicles from non-human intelligences (NHI) abound in popular culture and the media. For this article, we utilize the number of such claims to estimate the number of years expected before the occurrence of a “catastrophic disclosure,” a term which was defined at the 2023 Sol Foundation’s inaugural conference as the accidental disclosure of conclusive evidence of the existence of NHI, outside of the control of human institutions, such as governments and militaries. Here, we will consider one possible example of this as the crash of a piloted spacecraft (or, an ET probe) in the middle of a busy metropolis, such as a city square, e.g., New York’s Times Square. The distribution of humans on Earth’s surface, the population as a function of time, and the fraction of individuals owning smartphones, also versus time, are each taken into account as foundations for a rigorous statistical analysis. The author adopts a skeptical / agnostic approach and doesn’t claim NHI or ET are real but uses their analysis as an educational example of critical thinking and application of standard statistical distributions to an issue which captures the imagination of the public like almost no other issue. Making the extraordinary assumptions that sentient species other than humans exist, are capable of constructing vehicles for transportation, and are sufficiently fallible that their technology can malfunction, it is possible to quantify potential answers to the question of how long it will be before smartphone imagery or video proofs appear on the internet and become irrevocable through classification in the modern era. Results of simulations of numerous different potential scenarios, of varying degrees of pessimism / optimism, indicate that, if NHI/ET is/are real, catastrophic disclosure may actually happen accidentally rather soon, with the mean expected year being 2040 $\pm$ 20 under the default assumptions.
1 Introduction, and Literature Review
The concept of the existence of conscious, sentient, and intelligent races of non-humans has captured the collective imagination of humanity for centuries, or even for millennia. The issue of their exact natures, as well as the question of technological capabilities enabling visits to the Earth, are separable from the question of existence. UFOs (Unidentified Flying Objects) and UAP (Unidentified Aerospace / Anomalous Phenomenon/a) are often conflated with the notion of spacecraft that are being piloted by NHI (non-human intelligence) including and especially ET (extra-terrestrials), in spite of these terms, especially the latter newer one, referring only to an unknown phenomenon, or phenomena, which may include anomalous atmospheric effects which are naturally occurring but are simply not yet understood Szydagis et al. (2023). Such conflation happens for good reasons, however – the measured motions, specifically the high velocities and accelerations, of at least some small fraction of observed, but non-identifiable, aircraft Knuth et al. (2019). Since late 2017 especially, discussions about UFOs, as well as about aliens, have once again been thrust into the limelight, within the mainstream media Cooper et al. (2017). 2023 “whistleblower” claims reignited the fires.
The ET hypothesis is arguably a scientific, reasonable, and rational one. The exploration of it can easily be justified by the discoveries of many thousands of extra-solar planets (or, exoplanets), which probably constitute only a small sample from billions, or even trillions, with $O$(10%) potentially habitable according to the anthropocentric habilitity criterion based on the correct host-star separation to permit the presence of water in liquid form, and oxygen in the atmosphere, thus not even counting life not-as-we-know-it, nor exo-moons. A relatively recent example of a discovery of multiple possibly human-habitable worlds announced via a NASA press conference is the TRAPPIST-1 planetary system Gillon et al. (2017). While the relativistic time dilation plus the Lorentz length contraction would both work in favor of any high-speed travelers of course, the problems of fuel and propellant with sufficient thrust for long-term high acceleration, shielding against cosmic radiation and fatal Doppler blue-shifted starlight, and celestial navigation remain unresolved – at least by modern-day human beings. A multi-billion-year-old civilization might become capable of overcoming all these engineering (not physical) difficulties involved, but that is purely optimistic speculation, with no complete explanation for a“hide and seek” type behavior, so it is best to adopt a model-independent approach to the question of NHI origin(s) and travel abilities.
As a result, this paper will make no assumption in that regard – NHI, if assumed to be real and able to reach Earth, may be traditional aliens, come from other dimensions or universes, or be intelligent lifeforms that co-evolved next to humans and are thus also native to Earth. Furthermore, we do not discount human explanations, e.g. time-travelers or present-day breakaway civilizations, nor the mundane explanations of many sightings. All of the quite wild hypotheses with no solid empirical evidence as yet for any of them are covered very well in other sources, such as Sanderson and Childress (2005) and Puthoff (2022). The only (extraordinary) assumptions being made for this research article are:
  1. 1.
    NHI (or, a very advanced yet unknown group of humans possessing extraordinary vehicles) may actually exist.
  2. 2.
    Regardless of point of origin or motivations: they possess high-tech craft they operate on/near Earth’s surface.
  3. 3.
    They are not supernatural nor infallible, making accidents such as unplanned (i.e., crash) landings fully realistic.
Given the three simple points above we can now ask the question of when good preliminary evidence would be captured by ordinary civilians utilizing, for example, smartphone camera technology, through random chance, by considering an additional, fourth assumption, of an annual crash rate, which can be grounded by a review of the literature regarding UFO crash claims. Though we failed to find any scholarly papers from (external and blind) peer-reviewed, high-impact researched-focused journals in the mainstream scientific community for this particular sub-topic (nor many on UFOs in general due to the enduring stigma/taboo), there is at least Maristela (2023) to build on. Moreover, these 4 books proved to be very useful resources – Schmitt and Randle (1991); Randle and Schmitt (1994); Randle (1995, 2010).
At the Sol Foundation symposium organized by Prof. Garry Nolan of Stanford University and held on November 17–18, 2023, some speakers used the phrase “catastrophic disclosure” to describe this case, where the military-industrial complex is not the player driving disclosure, but scientists, engineers, and researchers, citizen scientists, and/or even ordinary citizens (with the unspoken postulate being that human governments know a great deal more than they are disclosing to date). Many of the Sol speakers suggested that in such a scenario the impacts upon society (on politics, religion, etc.) of disclosure of the existence of NHI would be more catastrophic if compared to a slow, controlled, and planned version of it. This article will present a skeptical, but not debunking, analysis, one that should allow readers, including relevant politicians and lawmakers, to estimate when catastrophic disclosure could transpire on its own, but also allow them to disprove the most extreme claims of high crash rates, especially as time goes on without disclosure transpiring. The statistical methods employed here may be germane to setting upper limits on the rate of occurrence of many different kinds of exotic phenomena. The particular case of a crashed NHI craft would most likely constitute the single most important discovery in the history of science, if not all human history in general.
2 Methods
To run Monte-Carlo computer simulations of random, smartphone-driven disclosure, three levels of crash rates were taken as benchmarks constituting simplistic, order-of-magnitude, Fermi-problem-style estimates – 1, 10, and 100 per century. The lowest value essentially comes from treating only the Roswell incident Schmitt and Randle (1991); Randle and Schmitt (1994); Birnes and Corso (2017) as a (potentially) real example of an NHI spacecraft crash, the sole one for the entire twentieth century, as suggested by K. Randle, who has stated that most other incidents were probably hoaxes and misidentifications. Our middle-of-the-road value of 10 stems from taking the claims of whistleblower David Grusch at face value Kean and Blumenthal (2023). The extremum of 100/century or 1/year originates from Randle (2010), which contains a list of 118, with most not NHI-related however, as already stated, thus making 100 not just the highest rate assumed within this work, but likely also the least realistic. That being said, Wood (2024) cites over 50 possibilities, using a rating scheme to judge the probabilities of their veracity, and Randle has written that lists of well over 300 alleged crashes exist. Therefore, 2 is a reasonable power of 10 for the upper end of our Fermi estimation.
Not only the planetary population as a function of time, projected into the future based on UN projections Raftery et al. (2014), but also the distribution of persons across the approximately 1.49 $\times$ 108 km2 of land Cohen and Small (1998), were taken into consideration. An average person density is not used to represent the entire globe. This would be unrealistic for areas such as Antarctica, Siberia, Greenland, and northern Canada at one extreme, and large, dense cities on the other (with Manhattan, NY not even being the densest example). The probability density function (PDF) for population density in units of people per km2 was modeled as a skew-Gaussian function, skewed in favor of lower densities, and spanning $10^{-2}$$10^{+5}$ individuals/km2. The PDF peaks at $\sim$10/km2, with an average of $\approx$30–60/km2, depending on the year being modeled. A robust density profile was required for this study, but the only scholarly one found was for 1998 Cohen and Small (1998). It was adjusted for later years by smoothly varying just the centroid of the skew-normal distribution to approximate new density distributions with time. This approximation, with all areas effectively going up in person density uniformly, was validated by integrating under the resulting curves and verifying that one recovers the correct total populations. Future work should account for the width and skew of the probability density changing over time as well, although adjusting these would be overkill for a zeroth-order analysis. (Note that, for reproducibility, all exact equations and numbers can be found in the appendix with the C code.)
The fraction of people who own smartphones (with cameras) versus time was also necessary to model; however, for simplicity we ignored any bias toward greater ownership in higher-population-density areas, applying only a flat value. That said, a Poissonian distribution, the most common assumption in STEM for the modeling of rare events, was used to simulate local variation in density of phone owners, with a Poisson random number generator likewise implemented to simulate the number of crashes (on land) each year (2008 and later), with rate estimates discussed earlier used only to set the Poisson means. Even though the Poisson function can be well approximated with a Gaussian or “normal” distribution (bell curve) at high rates, it has the advantages over that of producing only non-negative, integer values, and because its mean and variance are equal, a separate value for the function’s width is not necessary to justify, unlike with Gaussians. While the first “smartphone” was invented in the early 1990s by IBM, and the first camera-phones in the late 90s or 2000, no year earlier than 2008 was considered in our analyses, a year after the introduction of Apple’s iPhone, which led to a greater explosion in ownership, with competing companies also making phones with cameras.
The mathematical formulae applied within the analyses presented herein, while detailed in the appendix, are also summarized here, and demonstrated visually in Fig. 1. The world’s population density profile probability density was:
$e^{\frac{-(x-\xi)^{2}}{2\times 1.711^{2}}}[1+\mathrm{erf}(-2.4\frac{x-\xi}{1.711\sqrt{2}})];\leavevmode\nobreak\ \mathrm{x\leavevmode\nobreak\ is\leavevmode\nobreak\ log_{10}(density).}\leavevmode\nobreak\ \xi=1.5467+4.2773\cdot 10^{-11}P;\leavevmode\nobreak\ \mathrm{P\leavevmode\nobreak\ is\leavevmode\nobreak\ overall\leavevmode\nobreak\ pop.}$ (1)
[…]
with three $P$-increase scenarios indicated from left to right (low, moderate, and high) and $t$ the time (in years). The base-10 log of the camera smartphone ownership fraction (dimensionless) vs. $t$ was modeled as an asymptotic S-curve:
$-0.26506+\frac{-1.5883+0.26506}{1+[\frac{(t-1999.5)}{12.387}]^{6.7706}}\leavevmode\nobreak\ \mathrm{where,\leavevmode\nobreak\ \textit{e.g.},\leavevmode\nobreak\ -0.26506\leavevmode\nobreak\ means\leavevmode\nobreak\ 10^{-0.26506}=54.3\%\leavevmode\nobreak\ phone\leavevmode\nobreak\ ownership.}$ (3)
Lastly, a radius defined a circle (approximated as flat) on the surface of the Earth, within which at least one individual with a phone is located, based on a Poisson-varied number of people, based on a mean density drawn from a skew-Gauss distribution (Eqn. 1). The default used was 0.150 km, dubbed the Powell Radius, because of the crude approximation mentioned in Powell (2024) of 500 ft. for reliable eyewitness testimony (sans images), but for the sky, not the ground.
Figure 1: A summary of all of the numerical inputs to the results. (a.) An amplitude-normalized PDF for the population density at several example years. (When the PDF is normalized by area instead, its peak is at 1.56 not 1.00 on the y-axis.) Both plot axes are logarithmic. Because the x-axis is cut-off at -2, i.e. 0.01 people/km2, there is a slight overestimate of total populations in the earliest years, although this is irrelevant since they are too early. A slight underestimate in later years was produced by cutting x off at +5, or 100,000 people/km2, which exceeds Macao (in China) and Monaco’s present-day values, of about 20,000 people/km2 (higher densities may be reasonable for the future, even if rare). (b.) The 3 United Nations projections applied to the non-normalized versions of the curves in the left graph, which uses only “medium” as an example, to set the proper totals. These correspond to the 3 formulae in Eqn. (2), a symmetric Gaussian (which unphysically collapses to 0 pop, but not for many centuries), an asymmetric sigmoid which asymptotes to just over 10 billion, and a quadratic polynomial, for the low, medium, and high cases, respectively. (c.) Actual data on smartphone ownership percentage as a solid red line Sui et al. (2021), with a symmetric-sigmoid curve fit to the log of the fraction in pink dots. When the data are plotted using a log-y scale, an inflection point which may be attributable to the introduction of the Apple iPhone is evident. Our sigmoidal fit unphysically asymptotes to 2.6% at left, but the fit is only applied to the post-2007 time period. (d.) The inset displays 3 possible futures: a value barely above 50% persists (pink dotted line again) and (linear) increases to 100% at 2 different slopes – a dark orange long-dashed line and light orange short-dashed line, the latter unrealistically reaching 100% in the late 2060s, but this is simply one, extremely optimistic, scenario, but justified based on publicly-claimed slopes Jejdling (2024); Laricchia (2024) (e.) Differing radii used as low, medium, and high scenarios to define minimum distances within which crashes are obvious and capturable. No thought is given to matters like object size and camera focal length, but a factor-2 uncertainty on radius should cover most such issues.
The Poissonian probability $p$ of 1 crash can be approximated as the rate $r$ via Taylor expansion, but probability $p$ [^1] The very same form as Eqn. (4) can model the $k$ people who see and successfully record an incident, replacing $r$ with the mean number of people $N$=$10^{x}\pi R^{2}$ multiplied by the camera-phone ownership fraction of Eqn. (3). ($x$ is the log density, and (people/km2)$\times$×km2=people) of $k$ crashes in a given year is more fully and correctly expressed as:
$p=\frac{\langle r\rangle^{k}}{k!}e^{-\langle r\rangle},\leavevmode\nobreak\ \mathrm{where\leavevmode\nobreak\ \langle r\rangle\leavevmode\nobreak\ is\leavevmode\nobreak\ the\leavevmode\nobreak\ expectation\leavevmode\nobreak\ value\leavevmode\nobreak\ for\leavevmode\nobreak\ the\leavevmode\nobreak\ annual\leavevmode\nobreak\ crash\leavevmode\nobreak\ rate,\leavevmode\nobreak\ and\leavevmode\nobreak\ k=0\leavevmode\nobreak\ or\leavevmode\nobreak\ a\leavevmode\nobreak\ positive\leavevmode\nobreak\ integer.}$ (4)
For instance, for the rate of 10/century or $\langle r\rangle$ = 0.1/year, the $p$ of $k$=0 is highest, at 0.90 or 90%, with a 9.0% chance of $k$=1, and 0.45% for $k$=2, and so forth. The virtual dice can be re-rolled once a day, week, month, or year in code, by adjusting the units on $r$, with no substantive change in the ultimate results. We opted for the annual timeframe for computational speed, and performed 100,000 trials for each of our 18 scenarios based on 9 distinct combinations of the low, medium, and high possibilities presented in this section, and two different starting years, 2008 and 2024, for initializing the simulations. The large number of trials (105) ensured that statistical uncertainties would be negligible for our final results, dominated instead by the systematic uncertainties of our choices of the quantitative assumptions. A 2008 start was used to validate our work, checking the probabilities that disclosure should already have happened.
3 Results
[…]
We are, however, assuming for all of our simulation results that captured evidence is qualitatively conclusive, without attempting to quantify what is meant by “conclusive” evidence in terms of video and/or photographic quality, duration or number of pictures, etc. We also postulate that one eyewitness with a phone is sufficient, due to the fact that they can rapidly share data with friends and family via e-mail and by text attachments, and on public-facing video-sharing websites such as YouTube, TikTok, or Vimeo, before the military or other authorities can rush to the scene to remove all physical evidence, then classify the incident. (Such actions have naturally been alleged by UFO conspiracy theorists going back decades.) Furthermore, having one initial witness should be sufficient due to one’s ability to call to others in high-population-density areas and text friends within areas of any density (this criterion can be easily modified in the code, however, and the effect on the result explored). While humans have always been social creatures, there was no texting nor camera-phones in pockets in Roswell, NM in 1947, nor in Kecksburg, PA in 1965, for instance.
Figure 2 presents all the sim results, divided by annual probability and cumulative (integrated) probability vs. time, and also separated by start year, 2008 or 2024, with cases having the latter as the input essentially taking the current state of affairs (lack of proof on the web) as a (Bayesian) prior. In those latter situations, over-optimistic postulates can artificially generate enormous immediate peaks in probability, for 2024–2025. Using our “central” postulates, the predictions are 2038 $\pm$ 24 (2008 sim start) and 2049 $\pm$ 23 (2024 start) for the year of the initial but “incontrovertible” evidence being shared on the internet (assuming survival of strict checks of fakery). The uncertainties $i.e.$ errors quoted here are non-Gaussian – they are simply raw standard deviations $\sigma$ of asymmetric data. The mean expectations are not the same as the median, RMS (root-mean-square) or mode (peak in probability), with that last parameter having no value higher than 2068 even when others extend into the 22nd century, due $e.g.$ to a global-pop peak.
Figure 2: Odds of catastrophic disclosure by year at left (a,c) on log axes, and summed at right (b,d), beginning in 2008 (top), and 2024 (bot). Pessimistic refers to all inputs “low” (Fig. 1), realistic medium, optimistic high. Roswell means 0.01 crash/year, Grusch 0.1/yr, and Randle 1, but * indicates he thinks that is far too high. Some stats are tabulated at the right.
4 Discussion and Conclusion
If NHI are real then the correct question to ask is not IF disclosure can be forced, but WHEN. Figure 2 right shows the probability always asymptoting to 100% eventually – when inputting the defaults, the chances (black curves) are 14–42% by 2027 and 39–59% by 2036. (Two values are quoted for each example year because of the differing beginning years.) This is because of the remarkable interconnectedness of the modern world, with cellular phones in individuals’ pockets and bags nearly everywhere even in developing nations, even when one makes the conservative claim of only a fixed 54% maximum for the percentage of human beings with camera-enabled smartphones. We did not account for people who have multiple phones and subscriptions, but this could be the cause of the significant tension between the sources we cited earlier (50 as opposed to about 75%), and this is why it is so important that divergent scenarios were studied. A further refinement to our study that readers can implement on their own since the code has been provided at the end is a mixing of low, medium, and high inputs instead of their simultaneous railing, which was done here for simplicity. Future work building upon this paper by this author and/or others could also account for daylight versus night-time hours, terrain type, attention spans, and other such factors, but the analyses presented herein represent a first stab, the first in the UAP literature as far as the author knows, at least in terms of serious (journal) publications.
An unspoken assumption baked into every analysis within this manuscript is that we are referring to unmistakable, clear photos of crashes on land only, potentially less ephemeral than UFOs in the sky, including if not especially over waters Sanderson and Childress (2005); Dennett (2018), but lands account for only 29.2% of the surface area of Earth. Therefore, $\langle r\rangle$ = 0.1/year (our middle-benchmark crash rate) becomes $>$ 0.3/year for the entire globe, with $>$ 0.2/year for only the water-covered regions. If one equates any NHI with ETs, and takes the estimated time frame of the first discovery of Earth by a space-faring race capable of interstellar travel as $\sim$1 million years ago Knuth (2024), then this implies over 200,000 defunct extraterrestrial vessels sitting on the bottom of the Pacific Ocean, and other oceans, and seas, lakes, and rivers. This is a very surprising number which motivates further searches like that described in Loeb et al. (2024), and this might even be an underestimate – one species may inform others of Earth’s and of humanity’s presence, possibly leading to a time-dependent $r$-value (increasing?) driven by the scientific curiosity of others.
Even though this work may read like it is a warning to government officials with knowledge of NHI or ET (again– should they really exist) that their time grows short to maintain full control of the narrative, it should not just sound alarm bells. The mathematics is neutral, in the sense that one could also capitalize on the procedures contained here to debunk NHI and UAP crash retrieval claims, as more years without disclosure pass by. The statistical techniques employed in this article can and should also be applied to cryptids, orbs, Earth lights, ball lightning, and any sort of “paranormal” claim that presently can boast of no proof that has convinced the majority of the scientific community, nor the public at large. The author was inspired by working on dark matter particles, a mainstream scientific topic with an enormous number of indirect i.e. observational clues coming from cosmology and astrophysics, but no conclusive, direct evidence in the laboratory as yet – one can set limits on the probability of their interaction with normal matter composed of atoms Aalbers et al. (2023). A scientist can set a limit on the rate of occurrence of any rare phenomenon using similar techniques, but without necessarily ruling it out completely. Not all exotic claims are false, as our history has repeatedly shown (atoms, germs, continental drift, meteors, air/space flight, nuclear power/bombs, relativity,…)
That being said, it would be a mistake to state all initially “crazy” ideas have eventually been proven true within the history of science, as such a claim would be far from accurate. However, subjective experiences do not constitute unambiguous final evidence in the physical sciences, making the raw data sets of well-calibrated scientific instruments absolutely critical to possess, not just information from the “human sensor,” witnesses with potentially faulty memories (adequate for legal systems and useful in the humanities and the social sciences, but insufficient in a physical science). While a smartphone is not such an instrument, it can collect what Prof. Garry Nolan referred to at the Sol Foundation inaugural meeting as “pre-data,” a step above anecdotal non-data at least, which can still be used for corroborating scientific data, and justify choices of sensors. But, without funding and publications in mainstream journals, progress will still be difficult. A “smoking gun” phone pic could precipitate increases in both of those, but would still not be a substitute for the study of crash parts in person, to look for evidence of NHI technology, such as advanced unknown alloys, or isotopic concentrations inconsistent with our solar system Nolan et al. (2022). The latter could be observed using mass spectrometry, or non-destructively using NAA (neutron activation analysis) Laine et al. (2023). While an initially agnostic approach to data-taking is important Szydagis et al. (2023) scientists must be permitted, sans fear of reputation loss, to entertain “exotic” hypotheses (NHI, ET) and consider if data do favor them Villarroel (2023).
Acknowledgements
The author acknowledges Prof. Eric Davis as well as David Grusch for inspiring this paper, and Mick West for his skepticism, and xkcd 1235. He thanks Prof. Kevin Knuth for first getting him interested in UAPs, and Gary Voorhis, leader of UAPx. He also thanks Prof. Brenda Denzler and Prof. Beatriz Villarroel for crucial criticisms of Version 1.
Appendix A Supporting Code
The C++ file catDisc.cpp is attached for the benefit of the technically-minded reader with knowledge of computer programming. It can, for instance, be easily compiled on Unix with the terminal command “g++ -Ofast catDisc.cpp -o catDisc.out” assuming that one has the g++ compiler installed (the -Ofast optimization flag is optional).
Appendix B Supporting Plots
This appendix contains an additional figure, to explore a more fine-grained variation of the crash rate, as opposed to just looking at three orders of magnitude, and of $R$.
Figure 3: The year of accidental disclosure with uncertainty as a function of $\langle r\rangle$, with all other numerical assumptions held fixed at their “medium” levels, for human population growth, smartphone penetration into the population, and visibility radius (in black: $R$ is varied in gray). The inset is a zoom-in for 10–100 crashes/century. $r>20$ is disfavored by sims at the Powell Radius (0.15 km). Discrete sim results are well fit continuously using y = 2008 + $m_{1}/r^{m_{2}}+m_{3}/r^{m_{4}}$, with both powers floating point.